OpenStudy (anonymous):

If 4−kx−x^2 is divided by (x−4) , the remainder is -4. Find k.

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

help me with this one @Loser66

OpenStudy (loser66):

did I owe you something in previous life??

OpenStudy (anonymous):

nooooo

OpenStudy (loser66):

no, just today and at most tomorrow, my class starts on Thursday, no more help

OpenStudy (anonymous):

Oh I better get in all my questions then :D

OpenStudy (loser66):

replace x = 4 and let it = -4 to find k

OpenStudy (anonymous):

ok let me try

OpenStudy (anonymous):

k=3?

OpenStudy (loser66):

nope.

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

I give up

OpenStudy (loser66):

replace x =4 first

OpenStudy (anonymous):

4??

OpenStudy (anonymous):

I thought you said-4

OpenStudy (loser66):

4 - k x - x^2 = -4 replace x = 4 4 -k *4 - (-4)^2 = -4 4k = 8-16= -8 k = -2

OpenStudy (loser66):

go to eat something, don't tag me from now.

OpenStudy (anonymous):

:(